OpenStudy (anonymous):

Please help with my converting polar equations to rectangular?? **medal and fan** 1. r = 1 2. r = 4 / 2 - 2sin θ 3. r = 3 / 2 - 6sin θ 4. r = 2 / 2 - cos θ

OpenStudy (anonymous):

Somebody please help? I don't know how to do this :(

OpenStudy (anonymous):

\[r = \sqrt{x^2+y^2}\] \[x=r \cos \theta\] \[y = r \sin \theta\]

OpenStudy (anonymous):

\[1. \sqrt{x^2+y^2}=1\] 2-4. Multiply by the denominator, distribute, then make the substitution